# GRE Suite: Geo-localization Inference via Fine-Tuned Vision-Language Models and Enhanced Reasoning Chains

### v3

@@ 2 Related Work
Image Geo-localization. Image Geo-localization is an important task in computer vision [75, 76, 74], spatial data mining [65], and GeoAI [66]. As shown in  Fig.˜2, previous work in image geo-localization can be divided into four main modes: classification mode, retrieval mode, extra information mode and RAG mode. (1) Retrieval mode treat the image geo-localization task as a retrieval problem, typically maintaining a database of images [52, 26, 72, 54, 73, 47] or a gallery of GPS coordinates [48]. They take the most similar images and GPS coordinates to the query image as the predicted values. However, maintaining a global-level image database or GPS gallery is infeasible. (2) Classification mode [38, 49, 29, 51, 34, 8] divide the entire earth into multiple grid cells and assign the center coordinates as predicted values. Models are then trained to classify the input image into the correct cell. However, if the actual location of the image is far from the center of the predicted cell, there can still be significant errors, even if the cell prediction is correct. (3) Extra information mode approaches [48] incorporate higher-level geographical information, such as continental-scale priors, to enhance performance. Nevertheless, this approach essentially provides partial solutions, contradicting the fundamental purpose of the task. (4) RAG mode [70, 18] leverage large language models by retrieving relevant image-GPS pairs as references to optimize predictions. While there are also some tries based on diffusion method like Around the world [10]. However, these approaches rely on establishing large-scale aligned databases. In contrast to existing global image geo-localization approaches, we propose a reasoning-based methodology that leverages both explicit and implicit geographical indicators within images to predict open-ended coordinate prediction. Recent advances in MLLMs have enabled novel approaches leveraging their reasoning capabilities for geographic inference. While some works  [9, 46, 24]employ explicit reasoning chains, they lack systematic evaluation of reasoning quality. Complementary work has developed datasets  [2, 41] and reinforcement learning frameworks  [46] to enhance human-like geospatial reasoning.
@@
Reinforcement Learning. Recently, with the emergence of reasoning models like OpenAI’s o1 [17] and Deepseek-R1 [11], the research focus in Large Language Models (LLMs) has increasingly shifted towards enhancing the models’ reasoning capabilities through reinforcement learning (RL) techniques. Studies have explored improving LLMs’ performance in reasoning tasks such as solving mathematical problems [39, 53, 57, 4, 28] and coding [15, 20, 64, 61]. A notable breakthrough in this area is Deepseek-R1-Zero [11], which introduced a new approach to achieving robust reasoning capabilities using RL merely, eliminating the supervised fine-tuning (SFT) stage. However, current research on RL-based reasoning has largely been confined to the language domain, with limited exploration of its application in multi-modal settings. For LVLMs, RL has primarily been used for tasks like mitigating hallucinations and aligning models with human preference [67, 43, 58, 27, 59, 69], but there remains a significant gap in research focusing on enhancing reasoning and visual perception of Large Vision Language Models. To address this gap, our work uses a novel reinforcement fine-tuning strategy , applying verifiable rewards with GRPO-based [39] RL to visual geo-localization tasks. Our approach aims to improve the performance of LVLMs in processing various geo-localization tasks, especially when the high-quality fine-tuning data is limited.
@@ 3.1 GRE30K Construction
In this section, we present GRE30K, a geo-localization reasoning dataset designed to enhance the visual reasoning capability of MLLMs. Specifically, GRE30K consists of GRE30K-CoT for cold-start Initialization and GRE30K-Judge for reinforcement learning. Examples of the generated data are provided in  Section˜A.1.
@@ 5 Experiment
Implementation details: We adopt Qwen2.5-VL-7B as base model, the SFT experiments are conducted with a batch size of 128, a learning rate of 1e-5, and training over 1 epochs. Then, we perform RL on our dataset and experiment with training subsets of 10k for a single epoch each. All experiments are conducted with Pytorch and 16 NVIDIA H20(96G) GPUs.
@@ 5.1 Comparison with State-of-the-art methods
Moreover, our approach exhibits a large gain on the more challenging GWS15k dataset, surpassing the previous SOTA model with significant accuracy improvements of +0.2%, +1.0%, +2.0%, +9.1%, and +4.2% in the 1km, 25km, 200km, 750km, and 2500km thresholds respectively. Our model achieves superior performance over previous state-of-the-art approaches while utilizing merely 5% of the data, compared to their use of the complete MP-16 dataset. The GWS15k contains samples that are uniformly sampled across the Earth and are not biased towards any specific geographic location. Moreover, the images in this dataset have a large distribution shift compared to the training set, making the geo-localization task tough and challenging for brute-force alignment approaches. Our substantial improvement can be attributed to effective reasoning that leverages both explicit and implicit geographical indicators within images.
Table 2: We compare the performance of GRE with the state-of-the-art methods on (a) Im2GPS3k [12] and (b) GWS15k [8] datasets. Our method yields consistent gains across datasets and different distance thresholds. $\dagger$} denotes transformer-based models.
@@ 5.3 Ablation Study
To evaluate the effectiveness of our training data and training strategies, we compare the model’s performance under four distinct training strategies: (1) applying Cold-start Initialization on our dataset, (2) further optimizing the GRE-CI with RL stage I, (3) further optimizing the GRE-CI with RL stage II, and (4) further optimizing the GRE-CI with RL stage I and stage II. As illustrated in  Table˜4(b), the application of CI on our dataset significantly enhances the model’s performance on both the coarse-grained (e.g., country, continent) and fine-grained (e.g., city, street) localization performance. For (2) and (3) , (3) reach a comparable performance and (2) dropped at some levels of granularity, attributed to the misalignment between training and test task (reward) types in Stage I. Overall, (4) demonstrates superior performance to (3) due to its more robust reasoning capabilities.
@@ C.1 More Details on Training
Please refer to  Fig.˜11 and  Fig.˜12 for more details.
@@ C.2 Additional Main Results
We also conduct evaluations on the Google StreetView dataset( Table˜6), where we observe similar performance trends. Additionally, we demonstrate the performance of our base model, Qwen-2.5VL series, on Im2GPS3k and GWS15k datasets( Table˜7(b)). The results align with our conclusions from the main results, further validating the effectiveness of our proposed training strategy.
@@
C.3 Qualitative Results
@@ D.1 Limitations
The primary limitations of GRE include (1) substantial computational resource requirements, specifically utilizing 8 NVIDIA H20 GPUs for model training, and (2) the associated API costs for dataset generation.

### v4

@@ 2 Related Work
Image Geo-localization. Image Geo-localization is an important task in computer vision [92, 93, 91], spatial data mining [82], and GeoAI [83]. As shown in  Fig.˜2, previous work in image geo-localization can be divided into four main modes: classification mode, retrieval mode, prior knowledge mode and RAG mode. (1) Retrieval mode treat the image geo-localization task as a retrieval problem, typically maintaining a database of images [61, 32, 89, 63, 90, 56] or a gallery of GPS coordinates [57]. They take the most similar images and GPS coordinates to the query image as the predicted values. However, maintaining a global-level image database or GPS gallery is infeasible. (2) Classification mode [44, 58, 35, 60, 40, 8] divide the entire earth into multiple grid cells and assign the center coordinates as predicted values. Models are then trained to classify the input image into the correct cell. However, if the actual location of the image is far from the center of the predicted cell, there can still be significant errors, even if the cell prediction is correct. (3) Prior Knowledge mode approaches [57] incorporate higher-level geographical information, such as continental-scale priors, to enhance performance. Nevertheless, this approach essentially provides partial solutions, contradicting the fundamental purpose of the task. (4) RAG mode [87, 21] leverage large language models by retrieving relevant image-GPS pairs as references to optimize predictions. While there are also some tries based on diffusion method like Around the World [11], with application of flow matching and diffusion [28, 78, 81, 80, 47, 16] . However, these approaches rely on establishing large-scale aligned databases. In contrast to existing global image geo-localization approaches, we propose a reasoning-based methodology that leverages both explicit and implicit geographical indicators within images to predict open-ended coordinate prediction. Recent advances in MLLMs have enabled novel approaches leveraging their reasoning capabilities for geographic inference. While some works  [10, 55, 27]employ explicit reasoning chains, they lack systematic evaluation of reasoning quality. Complementary work has developed datasets  [2, 49] and reinforcement learning frameworks  [55] to enhance human-like geospatial reasoning.
@@
Reinforcement Learning. Recently, with the emergence of reasoning models like OpenAI’s o1 [20] and Deepseek-R1 [12], the research focus in Large Language Models (LLMs) has increasingly shifted towards enhancing the models’ reasoning capabilities through reinforcement learning (RL) techniques. Studies have explored improving LLMs’ performance in reasoning tasks such as solving mathematical problems [45, 62, 70, 4, 34] and coding [18, 23, 79, 74]. A notable breakthrough in this area is Deepseek-R1-Zero [12], which introduced a new approach to achieving robust reasoning capabilities using RL merely, eliminating the supervised fine-tuning (SFT) stage. However, current research on RL-based reasoning has largely been confined to the language domain, with limited exploration of its application in multi-modal settings. For LVLMs, RL has primarily been used for tasks like mitigating hallucinations and aligning models with human preference [84, 71, 33, 86, 68, 77, 48, 69, 67]. Interpretable visual reasoning, once a longstanding challenge [14], now benefits from RL-finetuned LVLMs acting as decision agents [73]. Cutting-edge models like Kimi [53] demonstrate advanced capabilities, with research expanding beyond hallucination mitigation to core reasoning enhancement. However, there remains a significant gap in research focusing on enhancing reasoning and visual perception of Large Vision Language Models. To address this gap, our work uses a novel reinforcement fine-tuning strategy , applying verifiable rewards with GRPO-based [45] RL to visual geo-localization tasks. Our approach aims to improve the performance of LVLMs in processing various geo-localization tasks, especially when the high-quality fine-tuning data is limited.
@@ 3.1 GRE30K Construction
In this section, we present GRE30K, a geo-localization reasoning dataset designed to enhance the visual reasoning capability of MLLMs. Specifically, GRE30K consists of GRE30K-CoT for cold-start Initialization and GRE30K-Judge for reinforcement learning. Examples of the generated data are provided in  Section˜A.1. While GWS15k [8] reveals Im2GPS3k [13]’s non-uniform distribution (with landmark repetition risks), our geographic filtering ensures clean evaluation.
@@ 5 Experiment
Implementation details: We adopt Qwen2.5-VL-7B as base model, the SFT experiments are conducted with a batch size of 128, a learning rate of 1e-5, and training over 1 epochs. Then, we perform RL on our dataset and experiment with training subsets of 10k for a single epoch each. All experiments are conducted with PyTorch and 8 NVIDIA H20(96G) GPUs.
@@ 5.1 Comparison with State-of-the-art methods
Moreover, our approach exhibits a large gain on the more challenging GWS15k dataset, surpassing the previous SOTA model with significant accuracy improvements of +0.2%, +1.0%, +2.0%, and +4.2% in the 1km, 25km, 200km and 2500km thresholds respectively. Our model achieves superior performance over previous state-of-the-art approaches while utilizing merely 5% of the data, compared to their use of the complete MP-16 dataset. The GWS15k contains samples that are uniformly sampled across the Earth and are not biased towards any specific geographic location. Moreover, the images in this dataset have a large distribution shift compared to the training set, making the geo-localization task tough and challenging for brute-force alignment approaches. Our substantial improvement can be attributed to effective reasoning that leverages both explicit and implicit geographical indicators within images.
Geo-localization in Vision-Language Models (VLMs) indeed highlights their ability to integrate world knowledge for inference—an emergent capability developed during training. To provide a comprehensive comparison, we have benchmarked both LLaVA-1.5  [31] and Molmo-D-7B [9] on the Im2GPS3k dataset, which use open-source training data.
Table 2: We compare the performance of GRE with the state-of-the-art methods on (a) Im2GPS3k [13] and (b) GWS15k [8] datasets. Our method yields consistent gains across datasets and different distance thresholds. $\dagger$} denotes transformer-based models. The asterisk (*) signifies that for a direct comparison, GeoReasoner was prompted to output coordinates, which differs from its default city-name output format.
@@
GeoReasoner [27] 0.2 1.6 2.1 3.9 6.8
GeoReasoner [27] 9.9 33.8 46.1 65.3 80.3
SeekWorld  [55] 4.3 29.8 44.9 59.1 67.3
Qwen2.5-VL-7B [3] 3.2 16.6 28.0 42.1 53.0
LLaVA-v1.5-7B [30] 1.7 7.5 11.3 20.8 44.6
Molmo-D-7B [9] 2.1 9.8 19.6 36.3 55.7
@@ 5.3 Ablation Study
To evaluate the effectiveness of our training data and training strategies, we compare the model’s performance under four distinct training strategies: (1) applying Cold-start Initialization on our dataset, (2) further optimizing the GRE-CI with RL stage I, (3) further optimizing the GRE-CI with RL stage II, and (4) further optimizing the GRE-CI with RL stage I and stage II. As illustrated in  Table˜4(b), the application of CI on our dataset significantly enhances the model’s performance on both the coarse-grained (e.g., country, continent) and fine-grained (e.g., city, street) localization performance. For (2) and (3) , (3) reach a comparable performance and (2) dropped at some levels of granularity, attributed to the misalignment between training and test task (reward) types in Stage I. Overall, (4) demonstrates superior performance to (3) due to its more robust reasoning capabilities. We also conduct additional ablation study on larger scale model and other open source model in Appendix C.2, materials can be found in the repository.
@@ C.1 More Details on Training
Please refer to  Fig.˜11 and  Fig.˜12 for more details. During the training process, the threshold is continuously updated. If the model can stably maintain enough rewards at the current granularity level, the threshold is further refined to a finer granularity level.
@@ C.2 Additional Main Results
We also conduct evaluations on the Google StreetView dataset( Table˜6), where we observe similar performance trends. Additionally, we demonstrate the performance of our base model, Qwen-2.5VL series, on Im2GPS3k and GWS15k datasets( Table˜7(b)). The results align with our conclusions from the main results, further validating the effectiveness of our proposed training strategy. We also have compared our model on the OSV-5M [2] in Table 8, where our model emonstrates excellent performance. As Around the World demonstrates excellent performance, we conduct study on MP16 dataset, and materials can be found in the repository.
@@
Table 8: Results on the recent OSV-5M [2] dataset
Method Street City Region Country Continent dist
11 km $25$ km $200$ km $750$ km $2500$ km Average Distance
Qwen2.5VL-7B 1.0 1.9 4.8 19.0 43.1 4942
Molmo-D-7B 0.7 1.1 1.3 7.2 32.1 6172
LLaVA-V1.5-7B 0.1 0.2 0.7 5.0 21.9 6895
SeekWorld 1.0 1.3 7.0 27.6 51.3 4326
SC Retrieval - 19.9 45.8 73.4 - 1386
RFM $\mathcal{S}_{2}$ - 5.4 44.2 76.2 - 1069
Ours 5.7 9.7 35.57 72.53 91.11 1192
C.3 Additional Ablation Study
We also conduct additional ablation study on Qwen2.5VL-32B and LLaVA-v1.5-7B [64](Table  9(b)), where we observe similar performance trends. The results demonstrate the efficacy and broad applicability of the proposed method.
Table 9: More ablation study on (a) Im2GPS3k [13] and (b) GWS15k [8] datasets.
(a) Results on the Im2GPS3k [13] dataset
Method Street City Region Country Continent
11 km $25$ km $200$ km $750$ km $2500$ km
LLaVA-v1.5-7B 1.7 7.5 11.3 20.8 44.6
CI(LLaVA) 4.2 10.2 24.9 42.9 58.9
CI + II(LLaVA) 6.1 14.6 31.3 47.6 63.1
Qwen2.5-VL-32B 6.5 25.1 41.0 59.9 75.3
CI(Qwen) 8.1 31.4 46.5 69.7 81.1
CI + I(Qwen) 7.6 30.1 40.2 71.2 82.2
CI + I + II(Qwen) 12.3 36.6 59.3 78.3 88.6
(b) Results on the recent GWS15k [8] dataset
Method Street City Region Country Continent
11 km $25$ km $200$ km $750$ km $2500$ km
Qwen2.5-VL-32B 0.06 0.36 7.5 28.5 52.4
CI 0.51 2.8 15.1 43.4 68.1
CI + I 0.42 2.3 13.9 43.6 68.1
CI + I + II 0.97 4.9 20.1 57.0 81.3
C.4 Qualitative Results
@@ D.1 Limitations
The primary limitations of GRE include (1) substantial computational resource requirements, specifically utilizing 8 NVIDIA H20 GPUs for model training, and (2) the associated API costs for dataset generation. GeoCLIP requires 155.63 GFLOPs per inference. In comparison, our model requires 262.27 GFLOPs for the visual encoder and 24,117.47 GFLOPs for the language model, which corresponds to 13.0506 GFLOPs per token. All FLOPs are measured using the THOP package.